# Random Matrix Theory-guided sparse PCA for single-cell RNA-seq data

### v1

@@ II Results
Having estimated $A$ and $B$, we now use them to denoise the biwhitened data $X\leftarrow A^{-1/2}XB^{-1/2}$. Recall that our strategy to guide sparse PCA with RMT is as follows: apply any sparse PCA method to $X$ and select its sparsity parameter $\gamma$ so that the inferred subspace $\hat{Q}$ forms the angle with the outlier eigenspace $W$ predicted by RMT. Specifically, the mapping between outlier and signal eigenvalues states that each outlier eigenvalue $\lambda$ corresponds to a unique signal eigenvalue $\alpha$ satisfying $\alpha=-1/\underline{m}(\lambda)$, where $\underline{m}(\lambda)$ is the complementary Stieltjes transform of $\rho_{S}$ (see Section IV). Likewise, each outlier eigenvector is expected to have squared overlap $\|Qv\|^{2}_{2}=\alpha\psi^{\prime}(\alpha)/\psi(\alpha)$ where $\psi(\alpha)$ is the functional inverse of $1/\underline{m}(\lambda)$, and is known analytically (see Section IV). Based on this, we propose choosing $\gamma$ such that the following relation holds:
@@
where $q=p/n$. If sparse PCA could exactly recover $Q$, this condition would hold with equality. In practice, exact recovery is not possible, so $\gamma$ must be chosen such that $\mathrm{tr}(\hat{Q}W)$ remains close to this lower bound.
As an important side note, unlike Eq. 8, Eq. 7 is fully general and also applies when working with left-whitened data, $X\leftarrow A^{-1/2}X$, rather than biwhitened data. This opens the possibility of applying the criterion directly at the count level, provided a better estimator for the support of $\rho_{S}$ becomes available for left-whitened data.
@@
Thanks to its generality, this criterion can be applied to any sparse PCA algorithms for which the sparsity level is controlled with a single parameter $\gamma$. The methods we consider are: i) max-variance (Gpower) [16, 17], ii) dictionary-learning (sklearn) [21], iii) regression-based (AManPG) [19, 20]. We also developed a naive approach based on the FISTA algorithm, detailed in Alg. 2 to solve a maximum-variance formulation of sparse PCA inspired by the SCoTLASS algorithm [18, 42]. When un-penalized, all these algorithms solve a different, yet mathematically equivalent, formulation of PCA [17]. When penalizing with an $L_{1}$ norm the loading vectors, they each solve nonequivalent problems, and we do not expect them to provide the same solution.
@@
Figure 2: RMT-guided sparse PCA accurately reconstructs the principal subspace of the data. a, b. Noise reduction as a function of the penalty for five different sparse PCA algorithms. For AManPG we use two $L_{2}$ penalties in its elastic-net regularization: no penalty (red) and a penalty of $10^{4}$ (green). For our FISTA implementation we compare two orthogonalization methods: Gram–Schmidt (brown) and Löwdin (blue). In pink, we apply sklearn algorithm after whitening instead of biwhitening. c. Error reduction at $\gamma=0.6\gamma^{*}$ for all datasets using the top-performing methods, with an average reduction of $\sim 30\%$. The colors correspond to the legend in a. For all datasets we used $2000$ highly variable genes, and preprocessing before biwhitening/whitening and sparse PCA followed the same steps as in Fig. 1.
@@
We now assess the viability of this criterion on realistic single-cell RNA-seq datasets, using the same seven technologies as before. For each dataset, we subsample $3000$ cells and select the $2000$ most highly variable features. After standard preprocessing: library-size correction and by log-normalization followed by biwhitening, we apply our approach to each subsampled dataset. To isolate the effect of biwhitening, we also substitute the biwhitening step by gene-wise $z$-scoring before applying our RMT-guided sparse PCA approach. Since the original datasets contain $\geq 30000$ cells, we have access to three subspaces: (i) the inferred subspace $\hat{Q}$, (ii) the outlier eigenspace $W$, and (iii) the outlier eigenspace for the full dataset, $W_{\mathrm{full}}$. In the limit $n\to\infty$, the outlier eigenspace converges to the true signal eigenspace $Q$, so we use $W_{\mathrm{full}}\simeq Q$ as a proxy for the signal. With this insight, we evaluate efficacy by measuring how much closer $\hat{Q}$ is to $W_{\mathrm{full}}$ compared to $W$. We define noise reduction as the relative improvement over classical PCA:
@@
In Fig. 2a, for the Zheng2017 and Macosko2015 datasets (see Section IV) we report the noise reduction for each sparse PCA algorithm as a function of $\gamma/\gamma^{*}$, where $\gamma^{*}$ is the optimal sparsity parameter for which criterion Eq. 8 is exactly satisfied. The curves collapse satisfactorily across methods, underscoring the generality of criterion Eq. 8. We also observe a sharp performance drop for penalty parameters above $\gamma^{*}$, illustrating the critical role of the criterion: applying sparse PCA with an overestimated penalty invariably destroys the biological signal. As noted earlier, because the inferred eigenspace does not perfectly match the true signal subspace, the criterion is not exact; in practice, selecting $\gamma\simeq 0.6\gamma^{*}$ appears to work well across all algorithms. Finally, Fig. S7 and Fig. S8 show that these results extend to all seven datasets and remain robust to changes in the set of highly variable genes.
@@
We also find that applying sparse PCA after gene-wise $z$-scoring leads to a dramatic loss in performance, highlighting the necessity of the biwhitening procedure. For the AManPG-based approach, the best results are obtained with elastic net regularization and a large $L_{2}$ penalty. As shown in Fig. S9, taking the $L_{2}$ penalty to diverge yields the strongest performance. By contrast, applying AManPG with purely $L_{1}$ regularization consistently underperforms relative to other sparse PCA methods. We interpret this as evidence that enforcing sparsity on regression weights, as in AManPG, is fundamentally different from imposing sparsity on loading vectors, even though the unpenalized problems are equivalent—a distinction recently emphasized in [44].
Overall, we obtain the best results with sklearn-, FISTA- and AManPG-based sparse PCA, with a clear preference for the pipeline combining library-size normalization, log-normalization, and biwhitening. For $\gamma\simeq 0.6\gamma^{*}$, this setup achieves an average noise reduction of $30\%$ over all the datasets, as shown in Fig. 2b. Surprisingly, our naive FISTA implementation for sparse PCA, where principal components are orthogonalized at each iteration using Löwdin’s method, outperforms all other approaches. This unexpected success suggests that, despite its heuristic nature, this new sparse PCA method stands as a top competitor among existing approaches.
@@
Figure 3: RMT-guided sparse PCA provides state-of-the-art performance on cell type classification. Improvement over PCA in out-of-bag error for bagging of $30$ $k$-NN classifiers with network weights inversely proportional to the distance between points and $n=25$ neighbors. The methods are ranked from best (top) to worst (bottom), as measured across classifiers using constant weights, inverse-distance weights, UMAP weights, and $n=10,25,45$ neighbors. To further reduce the variance of the classification error, we exclude from the analysis cell types represented by fewer than $30$ cells. The best-performing method (grey) uses all cells in the dataset ($\geq 30000$) to compute PCs, which are then used to project the subset of cells used in the benchmark. The next best performers are all RMT-guided sparse PCA methods. We use $3000$ cells and $2000$ highly variable genes. Complete benchmark flow charts for each dataset are shown in Fig. S3-S5. Each experiment was repeated $10$ times with a different cell subset selected at random. Benchmark results for other types of $k$-NN classifiers are shown in Fig. S10.
@@
We now evaluate whether sparse PCA improves unsupervised cell annotation compared to classical PCA. For this, we use three datasets with ground-truth cell type labels: (i) Zheng2017, human Peripheral Blood Mononuclear Cells (PBMC) annotated by correlation-based assignment to purified cell types [37]; (ii) Stuart2019, human Bone Marrow Mononuclear Cells (BMMC) annotated with an unsupervised CITE-seq approach integrating protein and RNA modalities [40]; and (iii) Luecken2021, human BMMC annotated using an unsupervised method with cross-modality validation (Multiome) [35]. On this task of unsupervised cell type annotation, we compare our approach with several methods designed to denoise single-cell RNA-seq data: (i) autoencoder-based methods: scVI [45] and DCA [46], (ii) the diffusion-based method MAGIC [47], and (iii) PCA- and RMT-based methods: scLENS [13] and BiPCA [11].
@@
For all methods, after selecting $2000$ highly variable genes, we apply the denoising algorithm either to raw count data (scVI and DCA) or to library-size–corrected and log-normalized data (MAGIC). As a control for the effect of biwhitening alone, after library-size correction and log-normalization, we also apply biwhitening followed by PCA. Performance is measured relative to the canonical pipeline: library-size correction, log-normalization, and gene-wise $z$-scoring followed by PCA—using the same number of components as in the other methods. Flow charts describing the benchmark design and processing pipelines for each dataset provided in Fig. S3-S5.
Because cell type annotation often relies on community clustering with a $k$-Nearest-Neighbor ($k$-NN) graph, it is especially relevant to evaluate performance at the level of nearest-neighbor relationships [48]. We assess the quality of cell type classification by measuring the accuracy of $k$-NN classifiers trained on the ground-truth labels. Following [49], classification performance is quantified using Bootstrap Aggregation, with the out-of-bag error serving as the measure of accuracy [50].
@@
In Fig. 3, we report the relative change in out-of-bag error for each method compared to PCA. By this metric, classification with sparse PCs performs on par with classification using PCs computed from the full dataset via $W_{\mathrm{full}}$. The latter serves as a control, representing the best performance achievable with the same number of PCs but many more cells. To have a fair comparison, we performed hyperparameter searches over $50$ models for both DCA and scVI. Since autoencoders embed data into a lower-dimensional space distinct from the PCs, we evaluated cell type classification both on PCs derived from denoised data and directly in their intrinsic latent space.
@@
Regardless of the latent space considered, all autoencoder-based methods underperformed relative to baseline PCA. Similarly, MAGIC also underperformed compared to PCA for this task. These results are consistent with those in Fig. S11, where performance is assessed by measuring the average silhouette score of the ground-truth clustering on each low-dimensional embedding. Because scLENS and BiPCA are designed to operate on more than the $2000$ highly variable genes typically used, Fig. S13 verifies that our conclusions are robust to the number of genes included. Finally, Fig. S12 confirms that our results do not depend on the specific method used to select highly variable genes.
@@
Our results support the conclusion that our approach—biwhitening followed by RMT-guided sparse PCA—is more effective than autoencoder- and diffusion-based methods for cell type annotation. Unlike autoencoders, which require fitting probabilistic models with thousands of parameters, our method is almost parameter-free. That said, in its current form it is tailored specifically to cell type classification, whereas autoencoders can address a broader range of tasks, including direct denoising at the level of counts.
III Discussion
@@
We evaluated our approach on seven datasets spanning seven single-cell RNA-seq technologies and across four sparse PCA implementations. In every case, biwhitening followed by RMT-guided sparse PCA produced subspaces much closer to the signal eigenspace than PCA alone, achieving average noise reduction of up to $30\%$. These gains translated directly to downstream tasks: on three datasets with ground-truth cell type labels, our method reduced $k$-NN classification error compared to PCA and PCA-based approaches, while autoencoder-based methods (scVI, DCA) and the diffusion-based method (MAGIC) failed to improve upon the PCA baseline. Moreover, in terms of $k$-NN classification, sparse PCA approaches achieved performance comparable to that obtained with PCA applied to nearly ten times more cells. Put simply, from the perspective of the PCs, using RMT-guided sparse PCA is equivalent to increasing sample size by an order of magnitude.
The main limitation of our methodology is that, in the absence of a better estimator for the support of $\rho_{S}$, we are constrained to operate on biwhitened data, for which the support of $\rho_{S}$ is analytically known. This restriction limits the scope of our approach to unsupervised classification tasks. In particular, exact counts are only recoverable after reverting the biwhitening process, yet there is no mathematical guarantee that this procedure would yield any improvement. A better estimator for the support of $\rho_{S}$ would overcome this limitation, enabling direct correction of gene expression at the level of counts by operating on left-whitened data.
IV Materials and Methods
Eigenspectrum of the sample covariance matrix

### v2

@@ 2.3 RMT-guided sparse PCA estimates the signal eigenspace better than standard PCA
Having estimated $A$ and $B$, we now use them to denoise the biwhitened data $X_{\mathrm{bw}}$. Recall that our strategy to guide sparse PCA with RMT is as follows: apply any sparse PCA method to $X_{\mathrm{bw}}$ and select its sparsity parameter $\gamma$ so that the inferred subspace $\hat{Q}$ forms the angle predicted by RMT with the outlier eigenspace $W$. Specifically, the mapping between outlier and signal eigenvalues states that each outlier eigenvalue $\lambda$ corresponds to a unique signal eigenvalue $\alpha$ satisfying $\alpha=-1/\underline{m}(\lambda)$, where $\underline{m}(\lambda)$ is the complementary Stieltjes transform of $\rho_{S}$ (see Section 4). Likewise, each outlier eigenvector is expected to have squared overlap $\|Qv\|^{2}_{2}=\alpha\psi^{\prime}(\alpha)/\psi(\alpha)$ where $\psi(\alpha)$ is the functional inverse of $1/\underline{m}(\lambda)$, and is known analytically (see Section 4). Based on this, we propose choosing $\gamma$ such that the following relation holds:
@@
where $q=p/n$. If sparse PCA could exactly recover $Q$, this condition would hold with equality. In practice, exact recovery is not possible, so $\gamma$ must be chosen such that $\mathrm{tr}(\hat{Q}W)$ remains close to this lower bound. As an important side note, unlike Eq. 8 which uses the analytical expression for $\psi(\alpha)$ for biwhitened data, Eq. 7 is fully general and thus also applies when working with left-whitened data, $X_{\mathrm{lw}}=A^{-1/2}X$, instead of biwhitened data. This opens the possibility of applying the criterion directly at the level of counts, provided a better estimator for the support of $\rho_{S}$ allows the identification of signal from noise with left-whitened data.
@@
Thanks to its generality, this criterion can be applied to any sparse PCA algorithms for which the sparsity level is controlled with a single parameter $\gamma$. The methods we consider are: i) max-variance (Gpower) [26, 13], ii) dictionary-learning (sklearn) [22], iii) regression-based (AManPG) [54, 14]. We also developed a naive approach based on the FISTA algorithm, detailed in Alg. S1 to solve a maximum-variance formulation of sparse PCA inspired by the SCoTLASS algorithm [25, 6]. When un-penalized, all these algorithms solve a different, yet mathematically equivalent, formulation of PCA [13]. When penalizing with an $L_{1}$ norm the loading vectors, they each solve nonequivalent problems, and we do not expect them to provide the same solution.
@@
Figure 2: Noise reduction obtained with sparse PCA. a. Noise reduction as a function of $\gamma/\gamma^{*}$ for five different sparse PCA algorithms. For AManPG we use two $L_{2}$ penalties in its elastic-net regularization: no penalty (red) and a penalty of $10^{4}$ (green). For our FISTA implementation we compare two orthogonalization methods: Gram–Schmidt (brown) and Löwdin (blue). In pink, we apply sklearn algorithm after gene-wise z-scoring (whitening) instead of biwhitening. b. Error reduction at $\gamma=0.6\gamma^{*}$ for all datasets using the top-performing methods, with an average reduction of $\sim 30\%$. The colors correspond to the legend in a. For all datasets we used $2000$ highly variable genes, and preprocessing before biwhitening/whitening and sparse PCA followed the same steps as in Fig. 1.
@@
We now assess the viability of this criterion on realistic single-cell RNA-seq datasets, using the same seven technologies as before. For each dataset, we subsample $3000$ cells and select the $2000$ most highly variable features. After standard preprocessing: library-size correction and log-normalization followed by biwhitening, we apply our approach to each subsampled dataset. To isolate the effect of biwhitening, we also substitute the biwhitening step by gene-wise $z$-scoring before applying our RMT-guided sparse PCA approach. Since the original datasets contain $\geq 30000$ cells, we have access to three subspaces: (i) the inferred subspace $\hat{Q}$, (ii) the outlier eigenspace $W$, and (iii) the outlier eigenspace for the full dataset, $W_{\mathrm{full}}$. In the limit $n\to\infty$, the outlier eigenspace converges to the true signal eigenspace $Q$, so we use $W_{\mathrm{full}}\simeq Q$ as a proxy for the signal. With this insight, we evaluate the efficacy by measuring how much closer $\hat{Q}$ is to $W_{\mathrm{full}}$ compared to $W$. We define noise reduction as the relative improvement over standard PCA:
@@
In Fig. 2a, for the Zheng2017 and Macosko2015 datasets (see Section 4) we report the noise reduction for each sparse PCA algorithm as a function of $\gamma/\gamma^{*}$, where $\gamma^{*}$ is the optimal sparsity parameter for which criterion Eq. 8 is exactly satisfied. The curves collapse satisfactorily across methods, underscoring the generality of criterion Eq. 8. We also observe a sharp performance drop for penalty parameters above $\gamma^{*}$, illustrating the critical role of this criterion: applying sparse PCA with an overestimated penalty invariably destroys the biological signal. As noted earlier, because the inferred eigenspace does not perfectly match the true signal subspace, the theoretical criterion is not exact. In practice, selecting $\gamma\simeq 0.6\gamma^{*}$ appears to work well across all algorithms, and we suggest adopting this empirical criterion when applying sparse PCA. Importantly, while this empirical criterion yields near-optimal performance, we also observe in Fig. 2a that any $\gamma\lesssim\gamma^{*}$ improves performance over PCA. Finally, Fig. S7 and Fig. S8 show that these results extend to all seven datasets and remain robust to changes in the set of highly variable genes.
@@
We also find that applying sparse PCA after gene-wise $z$-scoring leads to a dramatic loss in performance, highlighting the necessity of the biwhitening procedure. For the AManPG-based approach, the best results are obtained with elastic net regularization and a large $L_{2}$ penalty. As shown in Fig. S9, taking the $L_{2}$ penalty to diverge yields the strongest performance. By contrast, applying AManPG with purely $L_{1}$ regularization consistently underperforms relative to other sparse PCA methods. We interpret this as evidence that enforcing sparsity on regression weights, as in AManPG, is fundamentally different from imposing sparsity on loading vectors, even though the unpenalized problems are equivalent, a distinction recently emphasized in [42]. Surprisingly, our naive FISTA implementation for sparse PCA, where principal components are orthogonalized at each iteration using Löwdin’s method [35], outperforms all other approaches. This unexpected success suggests that, despite its heuristic nature, this new sparse PCA method stands as a top competitor among existing approaches.
Overall, we obtain the best results with sklearn-, FISTA- and AManPG-based sparse PCA, with a clear preference for the pipeline combining library-size normalization, log-normalization, and biwhitening. For $\gamma\simeq 0.6\gamma^{*}$, this setup achieves an average noise reduction of $30\%$ over all the datasets, as shown in Fig. 2b. With this benchmark, we showed that under the separable covariance model, sparse PCA improves the recovery of the low-rank signal over PCA. This, however, doesn’t necessarily indicates that sparse PCA improves biological signal recovery. For this reason, we now turn to evaluate the performance of our approach on the task of cell type annotation.
@@
Figure 3: Comparison of cell type annotation performance for state-of-the-art methods. Improvement over PCA in out-of-bag error for bagging of $30$ $k$-NN classifiers with network weights inversely proportional to the distance between points and $n=25$ neighbors. The methods are ranked from best (top) to worst (bottom), as measured across classifiers using constant weights, inverse-distance weights, UMAP weights, and $n=10,25,45$ neighbors. To further reduce the variance of the classification error, we exclude from the analysis cell types represented by fewer than $30$ cells. The best-performing method (grey) uses all cells in the dataset ($\geq 30000$) to compute PCs, which are then used to project the subset of cells used in the benchmark. The next best performers are all RMT-guided sparse PCA methods. We use $3000$ cells and $2000$ highly variable genes. Complete benchmark flow charts for each dataset are shown in Fig. S3-S5. Each experiment was repeated $10$ times with a different cell subset selected at random. Benchmark results for other types of $k$-NN classifiers are shown in Fig. S10.
2.4 RMT-guided sparse PCA outperforms state-of-the-art methods on cell type annotation
@@
We use three datasets with ground-truth cell type labels: (i) Zheng2017, human Peripheral Blood Mononuclear Cells (PBMC) annotated by correlation-based assignment to purified cell types [53]; (ii) Stuart2019, Human Bone Marrow Mononuclear Cells (BMMC) CITE-seq data, annotated with an unsupervised approach integrating protein and RNA modalities [46] ; and (iii) Luecken2021, Human BMMC Multiome data, annotated using an unsupervised method with cross-modality validation [36]. On this task of cell type annotation, we compare our approach, biwhitening followed by sparse PCA (termed SPCA), with several methods designed to denoise single-cell RNA-seq data: (i) autoencoder-based methods: scVI [34] and DCA [17], (ii) the diffusion-based method MAGIC [48], and (iii) PCA- and RMT-based methods: scLENS [27] and BiPCA [45].
@@
For all methods, after selecting $2000$ highly variable genes, we apply the denoising algorithm either to raw count data (scVI and DCA) or to library-size–corrected and log-normalized data (MAGIC). As a control for the effect of biwhitening alone, after library-size correction and log-normalization, we also apply biwhitening followed by PCA. Performance is measured relative to the canonical pipeline, simply referred to as PCA: library-size correction, log-normalization, and gene-wise $z$-scoring followed by PCA, using the same number of components as in the other methods. Flow charts describing the benchmark design and processing pipelines for each dataset provided in Fig. S3-S5.
@@
Because cell type annotation often relies on community clustering with a $k$-Nearest-Neighbor ($k$-NN) graph, it is especially relevant to evaluate performance at the level of nearest-neighbor relationships [1]. We assess the quality of cell type classification by measuring the accuracy of $k$-NN classifiers trained on the ground-truth labels. Following [32], classification performance is quantified using Bootstrap Aggregation, with the out-of-bag error serving as the measure of accuracy [10]. In Fig. 3, we report the relative change in out-of-bag error for each method compared to PCA. By this metric, classification with sparse PCs performs on par with classification using PCs computed from the full dataset via $W_{\mathrm{full}}$. The latter serves as a control, representing the best performance achievable with the same number of PCs but many more cells. For a fair comparison, we performed hyperparameter searches over 50 models for both DCA and scVI, using their built-in parameter tuning tools. Since autoencoders embed data into a lower-dimensional space distinct from the PCs, we evaluated cell type classification both on PCs derived from denoised data and directly in their intrinsic latent space.
@@
Regardless of the latent space considered, all autoencoder-based methods underperformed relative to PCA. Similarly, MAGIC also underperformed compared to PCA for this task. These results are consistent with those in Fig. S11, where performance is assessed by measuring the average silhouette score of the ground-truth clustering on each low-dimensional embedding.We also notice that biwhitening followed by PCA leads to better performance than PCA alone, which we interpret as improved variance stabilization and better identification of the signal eigenspace. Importantly, we observe an overall underperformance of methods applied directly to counts (BiPCA and SPCA on counts) compared to PCA applied to log-normalized data. Put simply, this suggests that log-normalization improves the quality of PCA-derived low-dimensional embeddings. This is consistent with the observation that log-normalization followed by PCA still outperforms alternative methods [1]. Finally, we show that SPCA with an overestimated sparsity parameter, $\gamma=5\gamma^{*}$, leads to drastic underperformance, highlighting that tuning $\gamma\lesssim\gamma^{*}$ is critical to SPCA’s success. Because scLENS and BiPCA are designed to operate on more than the $2000$ highly variable genes typically used, Fig. S13 verifies that our conclusions are robust to the number of genes included, and Fig. S12 confirms that our results do not depend on the method used to select highly variable genes.
@@
Our results support the conclusion that our approach, biwhitening followed by RMT-guided sparse PCA, is more effective than autoencoder- and diffusion-based methods for cell type annotation. Unlike autoencoders, which require fitting probabilistic models with thousands of parameters, our method is almost parameter-free. That said, in its current form it is tailored specifically to cell type classification, whereas autoencoders can address a broader range of tasks, including direct denoising at the level of counts.
3 Discussion
@@
We evaluated our approach on seven datasets spanning seven single-cell RNA-seq technologies and across four sparse PCA implementations. In every case, biwhitening followed by RMT-guided sparse PCA produced subspaces much closer to the signal eigenspace than PCA alone, achieving average noise reduction of $\sim 30\%$. These gains translated directly to downstream tasks: on three datasets with ground-truth cell type labels our method reduced $k$-NN classification error compared to PCA and PCA-based approaches, while autoencoder-based methods (scVI, DCA) and the diffusion-based method (MAGIC) failed to improve upon the PCA baseline. Moreover, in terms of $k$-NN classification, sparse PCA approaches achieved performance comparable to that obtained with PCA applied to nearly ten times more cells. Put simply, from the perspective of the PCs, using RMT-guided sparse PCA is equivalent to increasing sample size by an order of magnitude.
The main limitation of our methodology is that, in the absence of a better estimator for the support of $\rho_{S}$, we are constrained to operate on biwhitened data, for which the support of $\rho_{S}$ is analytically known. This restriction limits the scope of our approach to providing low-dimensional embeddings with improved signal-to-noise ratio compared to those inferred via standard PCA. Starting from these denoised low-dimensional embeddings, however, it remains unclear how to denoise the raw data itself. While it is tempting, as suggested by our empirical results in Fig. 3 with the unwhitening version of BiPCA, to simply revert the biwhitening procedure, we are not aware of any mathematical guarantee that this will consistently improve the biological signal relative to the raw data. A better estimator for the support of $\rho_{S}$ would bypass this issue altogether, enabling us to identify, and subsequently correct, the signal eigenspace directly on raw data.
4 Methods
4.1 Eigenspectrum of the sample covariance matrix